carbide, COC 2 an important preliminary chemical for industries producing synthetic fabrics and plastics. COC 2 may be produced by heating calcium oxide with coke C 2 O+3C COC 2 +CO What is the amount of CaC 2 which can be produced from the reaction of excess calcium oxide and g of carbon (Assume 100% efficiency of reaction for purposes of this problem ) (Atomic mass C = 12 O=16,Cq=40\
1
Expert's answer
2020-11-10T14:06:29-0500

CaO + 3C → CaC2 + CO

m(C) = 10.2 g

n = m/M

"n(C) = \frac{10.2}{12} = 0.85 \;mol"

"n(CaC_2) = \frac{1}{3}n(C)"

"n(CaC_2) = \frac{1}{3} \times 0.85 = 0.283 \;mol"

M(CaC2) = 40 + 24 = 64 g/mol

"m(CaC_2) = 0.283 \times 64 = 18.11 \;g"

Answer: 18.11 g